OpenStudy (anonymous):

For the function f(x) = 2 – 3x, find f(–4). A. –10 B. –5 C. 12 D. 14 @MATTW20

OpenStudy (anonymous):

just plug in -4=x

OpenStudy (anonymous):

and solve

OpenStudy (anonymous):

-5?

OpenStudy (anonymous):

what is -4*-3?

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

no not plus but times

OpenStudy (anonymous):

I mean -4 * -3

OpenStudy (anonymous):

yes, sorry, maybe I wasn't clear

OpenStudy (anonymous):

So what is -4 * -3?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

and what is 2+12

OpenStudy (anonymous):

14

OpenStudy (anonymous):

correct. So do you understand what you're doing here? You just put the 4 where the x would be and solve

OpenStudy (anonymous):

4= 12-14

OpenStudy (anonymous):

??

OpenStudy (anonymous):

hmmmmmmmmm.... how to explain this better

OpenStudy (anonymous):

f(4)= 2 - 3x

OpenStudy (anonymous):

^^^

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so now where the x is you just put the 4 and solve so f(4)=2-3(4)

OpenStudy (anonymous):

and the same thing goes for any number. If i said solve for f(100) you would just put f(100)=2-3(100)

OpenStudy (anonymous):

and i meant -4 in that first one

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

for what?

OpenStudy (anonymous):

this question lol

OpenStudy (anonymous):

what did you get lol we got the answer a while ago I just wanted to make sure you understood the concept. It's 14 remember

OpenStudy (anonymous):

ohh we did well i got 14 lol

OpenStudy (anonymous):

yup congrats

OpenStudy (anonymous):

thanks lol

OpenStudy (anonymous):

you're very welcome